OpenStudy (anonymous):

Derek orders a 6-ft sub sandwich for himself and two friends. Cory wants his piece to be 2 feet longer than Tamara's piece, and Tamara wants half as much as Derek. Find the length of each person's sub. For some reason my head can't wrap around this question. I keep reading it and don't know where to start...

OpenStudy (anonymous):

tamara wants half dereks = 1/2 * 6 = 3 feet cory wants 2 feet longer than tamara = 3 + 2 = 5 feet

OpenStudy (plainntall):

Have to write it down. Let us say that Derek's piece is x long. That means that Tamara's is half of that or x/2 Cory is 2 feet longer than Tamara's which is (x/2)+2 All of that has to add up to 6 feet long so x +(x/2)+((x/2)+2)=6 can you solve that equation?

OpenStudy (plainntall):

The trick is to find a variable that all the other quantities can be expressed in and then solve for that variable.